# Turán number of complete multipartite graphs in multipartite graphs

### v1

@@ 1. Introduction
To give the precise value of $\operatorname{ex}_{k}(n,K_{r+1}(t))$, we need the following definitions. Given $a,t,n_{1},\dots,n_{a}\in\mathbb{N}$, let $z_{t}^{(a)}(n_{1},\dots,n_{a})$ be the $a$-partite Zarankiewicz number for $K_{t,t}$, namely, the maximum number of edges in a $K_{t,t}$-free $a$-partite graph with part sizes $n_{1},\dots,n_{a}$. We also write $z_{t}^{(a)}(n)$ for $z_{t}^{(a)}(n,\dots,n)$. Given $r,k,t,n\in\mathbb{N}$ with $k\geq r+1$, assume $k=ar+b$ for some $10\leq b\leq r-$ 1, in particular, $b\geq 1$ if $a=1$. Let
$t_{1}:=\left\lfloor\frac{t-1}{4a^{2}+6a}\right\rfloor,\quad\text{and}\quad b^{\prime}:=\min\left\{b-1,\left\lfloor\frac{r-b}{2}\right\rfloor\right\}.$
Note that $b^{\prime}=b$ if and only if $b\leq(r+2)/3$. The following two functions will appear in our lower and upper bounds for $\operatorname{ex}_{k}(n,K_{r+1}(t))$, respectively.
$h_{1}(n,r,k,t):=\begin{cases}\lfloor\frac{t-1}{2}(k-a-1)n\rfloor&\text{ if $a\geq 2$, $b\geq 2$ and $t_{1}=0$},\\ \lfloor\frac{t-1}{2}(k-a-1)n\rfloor+\min\{b-1,r-b\}\frac{(2a^{2}+3a)t_{1}^{2}}{2}-\frac{r-1}{2}&\text{ if $a\geq 2$, $b\geq 2$ and $t_{1}>0$},\\ (r-1)\lfloor\frac{t-1}{2}an\rfloor&\text{ if $a\geq 2$ and $b=0,1$},\\ (t-1)(b-1)n+b^{\prime}\lfloor\frac{(t-1)^{2}}{4}\rfloor&\text{ if $a=1$ and $b\geq 1$},\\ \end{cases}$
and
$h_{2}(n,r,k,t):=\begin{cases}\lfloor\frac{t-1}{2}(k-a-1)n\rfloor+(r-b)\frac{(t-1)^{2}}{16(a-1)}&\text{ if }a\geq 2,t\geq a,\\ \lfloor\frac{t-1}{2}(k-a-1)n\rfloor&\text{ if }a\geq 2,t<a,\\ (r-1)\lfloor\frac{t-1}{2}an\rfloor&\text{ if $a\geq 2$ and $b=0,1$},\\ (t-1)(b-1)n+(b-1)\lfloor\frac{(t-1)^{2}}{4}\rfloor&\text{ if $a=1$ and $b\geq 1$}.\end{cases}$
Note that $h_{1}(n,r,k,t)=h_{2}(n,r,k,t)$ in the following cases:
$(1)\ a\geq 2\text{ and }t<a,\quad(2)\ a\geq 2\text{ and }b=0,1,\quad(3)\ a=1\text{ and }1\leq b\leq(r+2)/3.$ (1.2)
For $i=1,2$, we accordingly define
$g_{i}(n,r,k,t):=t_{r}(k)n^{2}+z_{t}^{(\lceil k/r\rceil)}(n)+h_{i}(n,r,k,t),$
and note that $\lceil k/r\rceil$ is equal to $a+1$ if $b>0$ and $a$ otherwise.
We have the following lower bound for $\text{ex}_{k}(n,K_{r+1}(t))$.

### v2

@@ 3.1. A stability theorem
In the earlier version of this paper, we gave a self-contained proof of Theorems 5. Here we derive it from a stronger result of Chen, Lu and Yuan [CLY23]*Theorem 1.5, in which they provide more structural information. Their definition and result are more general by allowing the parts $V_{i}$ of $G$ to have different sizes. Here we only state their definition and result that we need for the balanced case.
Definition 3.2 (Stable partition).
Let $k\geq r\geq 2$ be integers. Let $\mathcal{P}=\{P_{1},\dots,P_{r}\}$ and $\mathcal{V}=\{V_{1},\dots,V_{k}\}$ where $|V_{i}|=n$ for all $i\in[k]$ be partitions of a common vertex set. For $i\in[r],j\in[k]$, a set $W=P_{i}\cap V_{j}$ is called an integral part if $W=V_{j}$, and called a partial part otherwise. We say that $\mathcal{P}$ is stable to $\mathcal{V}$, if each of $P_{1},\dots,P_{r}$ has the same number of integral parts and at most one partial part.
Definition 3.3 ($(X,\varepsilon)$-stable).
Let $n,r,k$ be integers with $n\geq k\geq r\geq 2$. For a given spanning subgraph $G$ of $K_{k}(n)$, let $\mathcal{P}=\{P_{1},\dots,P_{r}\}$ be a partition of $V(G)$ and $\mathcal{V}=\{V_{1},\dots,V_{k}\}$ be the natural $k$-partition. Given $\varepsilon\in(0,1)$ and a set $X\subseteq V(G)$ of size at most $\varepsilon n$, we say that $\mathcal{P}$ is an $(X,\varepsilon)$-stable partition if
  • $G-X$ is $\varepsilon$-close to $K(P_{1},\dots,P_{r})-X$,
  • $\{P_{1}\setminus X,\dots,P_{r}\setminus X\}$ is stable to $\{V_{1}\setminus X,\dots,V_{k}\setminus X\}$.